HUBBLE SPACE TELESCOPE DAILY REPORT #831 PERIOD COVERED: 0800 EST Daily Status Report as of 056/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of HSP/UV1 4209 (HSP Detector Sensitivity Test) The High Speed Photometer was used to make four ultraviolet observations of standard star VID998 to measure the sensitivity of its detector tubes. The observations were executed in fine lock. The data was displayed for a representative of the principal investigator. No problems were reported. 1.2 Completed Three Sets of WFC 4713 (WF/PC Calibration: WFC Internal Delta Flats) The Wide Field Camera was used to make one internal flat observation. This image will be used for the restoration of Earth Flats, which are altered due to the changes in quantum efficiency which occurred after WFPC decontamination operations. The image was received and displayed. The internal flat was reported to look normal. 1.3 Completed HSP/UV2 1095 (Variability of High Luminosity Stars) The High Speed Photometer was used to make three observations of one of the most luminous and massive stars known, HD93250. All three observations were made in fine lock. A recentering event occurred during the last observation. All of the images were displayed for the principal investigator. 1.4 Completed FGS 4064 (Crit. Test of Galactic Escape Velocities at R(Sun)) The FGSs were used in their astrometric mode to make positional measurements of star G1625 to determine its parallax and proper movement. The observations were made in fine lock. One recentering event was seen, and no other problems were reported. 1.5 Completed Two Sets of WFC 4105 (HST Medium Deep Survey: High Latitude) The Wide Field Camera was used to make a total of four medium-deep sky survey images taken in parallel with the prime science observations. All of the observations were taken in fine lock. A recentering event occurred during the first image. The other observations were reported to be normal. 1.6 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.7 Completed PC 3209 (PC Images of Eta Carnae Core) The Planetary Camera was used to make a series of 24 observations of the Star System Eta Carnae. The first series was received and where possible displayed. They were reported to look normal. The last ten observations still remain on the science tape recorder, and will be analyzed later. 1.8 Completed PC 3603 (Snapshots of Protoplanetary Nebulae) The Planetary Camera was used to make observations of the protoplanetary system CPD-53D573. The observations were executed as scheduled, and no problems have been reported. 1.9 Completed FOS/RD 4495 (FOS GIMP Test) The Faint Object Spectrograph was used to initiate a spectral test to measure geomagnetically induced movement in its detector tubes. (See Operations Note 2.4) The test is currently under way, and no problems have been reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 0 Successful : 0 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Significant HSTARs: None 2.4 Operations Note: A Geomagnetic Induced Movement Problem (GIMP) Test is currently under way. The Red Side of the test has been completed (056/0500Z-0930Z). The test results were displayed in the OSS, and the image motion has been taken out. This section of the test appears to be a success. The Blue Side of the test is still ongoing (056/11Z-1600Z). 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Three Sets of PC 3603 (Snapshot of Protoplanetary Nebulae) 3.2 Complete FOS/RD 4495 (FOS GIMP Test) 3.3 Complete FOC/96 3790 (Polarization Imaging of Radio Galaxies) 3.4 Complete GHRS/2 3614 ( boron Probe of Stellar Structure and Mass Loss) 3.5 Complete Two Sets of WFC 4713 (WF/PC Calibration: Internal; Delta Flats) 3.6 Complete WFC 3253 (Optical and UV Observations of Radio Pulsars) 3.7 Complete PC 4140 (WF/PC Calibration: Rapid Internal Monitor Part 2) 3.8 Complete Three Sets of PC 4714 ( WF/PC Calibration: PC Internal Delta Flats) and Initiate a Fourth 3.9 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.10 Complete HSP/POL 4209 (HSP Detector Sensitivity Test) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-